OpenStudy (anonymous):

Find the slope of the line that contains the points named. K(c + d, c - d), L(c - d, c + d) m = ? the formula for slope is y2-y1 over x2-x1= the slope

OpenStudy (anonymous):

m = -1 K(c + d, c - d), L(c - d, c + d) Using the slope formula, m = [(c+d)-(c-d)] / [(c-d)-(c+d)] = (c + d - c + d) / (c - d - c - d) = 2d / -2d = -1

OpenStudy (anonymous):

Thanks again... i see where i made my mistake. :-)